Initial Setup at Front Bent Gregorian - 16 September 2011 UT

Observer(s): J. Hill (Tucson)
Telescope Operator: S. Allanson (LBTO)
Telescope Support: K. Newton (LBTO)
Instrument Support: T. Shih (LBTO)
Software Support: N. Cushing (LBTO)

Summary

It was pretty good in the beginning, but then both the weather and technical things fell apart.

We opened at 04:30 with clearish skies.

A communication problem with the SX adaptive secondary was apparently caused and later cured by Norm doing something with fibers. Doug and Guido were also fighting some kind of IDL problem. See Issues 3628-3630.

We measured the pointing and collimation offsets for RFBG in a short period of time.

When we got an image with LFBG it had mammoth coma. Something like -30000 nm of Z8 would be needed to correct it (beyond the range of active correction on M2). Without worrying about pointing, it looks like translating SX M2 to X=+8.00 mm reduces the coma. A summary of my present worries are: Did we move the swing arm the wrong amount last week? Or were we working on some kind of ghost image tonight?

We closed with heavy clouds (and soon mist) at 09:15 UT before we reached a firm conclusion about the coma problems.

We had minor annoyances from MCSPU, AOS, OSS and GCS-AzCam. Shutter doors continue to require latching and unlatching in Manual Mode.

Plan for Friday Night

My plan is to spend the first half working on the SX coma problem, with the main emphais on making sure we are looking at the real image and not a ghost. Then handover to Olga for the second half (more stable) for LBC verification.

Preparations (the last thing we did to SX last week)

Date: Thu, 8 Sep 2011 18:27:15 -0700
From: John Hill 
To: telescopework@lbto.org
Subject: Adjusting SX M2 Swing Arm Position

We (Jeff, Geno, John, Doug) adjusted the SX M2 Swing Arm so that the
mirror is 7 mm farther inboard and 2+ mm farther +Y.  We loosened the
adjusting nuts by 3.5 flats = 7/8 turns = 3.5 mm.

Details

SX AO secondary is not able to turn on. Doug and Guido are working on it.

04:15 Trying to open shutter doors. They remain uncooperative in automatic mode. Inboard pins are not able to unlatch/latch automatically. Kevin had to reset the circuit breaker on Latch Pin 1001.

04:30 Shutter Doors are Open! Moving to zenith to let Norm fix a fiber issue he identified with AO secondary.

04:45 SX AO M2 is working now that the fibers are reconnected. Doug and Guido are still dueling with some IDL software problems.

As IRTC at RFBG (PISCES there but not operating)

We are working on RFBG for two reasons: a) SX M2 isn't working, and b) we want to use DX pointing to verify the mount position for SX.

*** Initial DX Pointing Model Terms ***  IRTC_RFBG_20101226.pdmod
IA      -359.43070
IE         4.71030
NPAE     -77.09970
CA         0.00000
AW         0.24250
AN        19.12040
HACE2    -51.28600
TF        18.05370

04:50 Preset to M5 (21:33:58 +45:35:29)

04:55 Position of DX M2 hexapod is not updating.

04:57 Kill and restart OSS. A star briefly flashed by while hexapod was updating (moving to 0,0,0,0,0).

04:59 ptspiral DX side - still DX M2 problems

05:00 Steve resends preset

05:02 ptspiral DX side - saw lots of faint stars, but failed to find the bright one.

05:15 Preset to Vega (alias M5_0699) - found at approx IE=-15 CA=-80, but too saturated to center.

05:20 Preset to ACT0305 IE=-10 CA=-84

05:27 backing out the IE/CA correction with DX M2 hexapod translation (without running any active optics)
DX M2 Xglobal is -3.8 mm to get CA back to the nominal 0.00. Xtotal=-1.145

05:29 RFBG rotator faulted. Go to zenith to reset .........also restarted MCSPU.

05:45 Called Dan Cox. (had to restart MCSPU a second time)

Installed the SXSMLUCIFERCollimation.dat file into SXSMIRTCFCollimation.dat file.

05:55 Preset to a higher ACT star........
DX M2 Xglobal=-3.9 Yglobal=-0.8 Zglobal=1 at EL=72

06:01 Active Preset to BS9188 at EL=75 - Seeing is 0.7-0.9 arcsec.
DX M2 Xactive=3.289 Yactive=-0.862 Zactive=-0.262 RX=78.3 RY=336.4
DX M2 Xtotal= 1.610 Ytotal=-4.497 Ztotal=1.920 RX=-46.4 RY=939.97 @ EL=76

Sum of the globals and active corrections to give the new global offsets:
X=-0.611 Y=-1.662 Z=+0.738 RX=78.3 RY=336.4

*We have completed the observations for Andrew's steps 1a) to 1e) on the RFBG side.*

JMH: DXSMIRTCFCollimation.dat was updated the following evening.

LFBG compared to RFBG

06:13 Authorize as IRTC now on both sides, and flatten SX M2. SX selector failed.

06:14 Init/Home SX M3

Thin clouds now covering entire sky.

06:18 Authorize as LUCI1 and IRTC - succeeded. Initialize SX M2 hexapod.

06:22 Wade calls to say that there is a major diesel leak in the utility bldg. We are now running the generator from the day tank which has 4 hours capacity max. (We don't know the margin under "max".)

06:31 Syncronized presets on left and right to M5_0835. Right side fails for OSS "Stop[20] complete".

06:35 Abort left preset and restart OSS.

LBFG alone

Home the SX M3 again, and authorize LUCI1@LFBG only.

06:40 preset on left to M5_0835 and ptspiral only found one ghost (or out of focus image).

Since we are LUCI1, this is a different pointing model than IRTC (either one). LUCIFERSX.20110213.ptmod

06:45 Mirror Ventilation stopped to conserve fuel.

06:55 Preset to Vega. Found a very comatic image near Vega that was greatly defocussed. IE=-35 CA=+43.

07:00 Preset to nearby ACTxxxx and we see something albeit faint.

07:05 Preset to ACT0336 star at higher elevation (found within 15 arcsec, still very comatic)
Backing out the IE/CA adjustments with X-Y translation (still very comatic, ~40 arcsec image).

*** Pointing Model Terms *** LUCIFERSX.20110213.ptmod
IA      -396.03370
IE       -72.9962
NPAE     -12.77660
AW         1.52190
AN        25.49120
CA        63.6732
HESE2     42.40920
TFP       20.09550
HSCA2     -2.48990
HECA3      1.44060
HSSA3     -1.28080

Xglobal went from +0.2 to -0.300, Yglobal went from 1.0 to 4.40, Zglobal went from -0.1 to 1.500 RX=-104 RY=+21

07:23 couldn't handle coma. Sent acquire preset ACT0336 but we lost the star. There are no clouds. All mirrors collimated.

07:29 ptspiral IE=-39 CA=+77 why did this change?

07:35 Making M2 RY more negative to reduce the massive coma (RYglobal now -560 arcsec removed vaguely half the coma RYtotal=)

07:40 Back out the pointing changes with M2 translations again. Xglobal=-4.7 Xtotal=-6.90 but now the coma is worse again.

07:49 Apply -5000 Z8 tp M2, then another -5000 Z8 (has now removed maybe 40% of the X-coma?) Xactive=+9.3 RY=+970

07:54 Preset to M5 star to check pointing

08:05 John Ratje and XXXXX have arrived to babysit and debug the diesel fuel issues, so we can continue observing a while longer.

Initialize the collimation and clear active optics. Now star is at:approx IE=-72 CA=+63

08:16 Z8 = -10000, then increasing to Z8=-13000 (net) which is as much as we can apply to M2.

08:24 Clear Active, Xglobal =+5.2 mm

08:26 Xglobal=10.0 which seems to have overshot the coma, but the result is confused and delayed by passing clouds.

08:36 Xglobal=7.5 then start a ptspiral, but too many clouds.

08:42 Authorize IRTC command on both sides hangs at "Set new instrument (GCS)" on both sides of IIFGUI.

08:52 Stop GCSL, Kill GCSR. Restart.

Close

09:00 Close for heavier clouds and some new radar echoes forming. Closing the doors still requires manual operation of the pins. Balls on 1003 would not extend.

09:10 GCSR still says "Recovering", so it seems as if the unknown AzCam problem (Issue 3610) has come back.

09:16 Stop both GCSs. Soft reboot azcamserver.

09:22 Restart both GCSs, and reauthorize IRTC on both sides. The authorization was successful.

09:25 flatten both shells

-- JohnHill - 16 Sep 2011
Topic revision: r4 - 17 Sep 2011, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback